Grade 12th passPhysical Chemistry

Question 01
FInd n if nP5=3n+1P4
Question 02
if nC7=nC11 find nC16

??? ????? ???????
9 Years agoGrade 12th pass

1 Answer

Arundhuti
9 Years ago
n!/(n-5)! = 3*(n+1)!/(n+1-4)!n!/(n-5)! = (3*(n+1)n!) / ((n-3)(n-4)(n-5)!)n^2 +12 -7n = 3n + 3n^2 -10n +9 = 0Solving the equationWe getn = 9 or 1As 1 cannot be the answerSo n = 9